OpenStudy (anonymous):

derivative of (x^2-4)/(x) I have 2x^2-x^2+4 so far...

OpenStudy (anonymous):

ohh divided by x^2

zepdrix (zepdrix):

Ohh yes :3 good catch hehe

OpenStudy (anonymous):

yes! lol but can it be reduced?:O

zepdrix (zepdrix):

One sec, I wanna check your work ^^\[\Large \left[\frac{x^2-4}{x}\right]'\quad=\quad \frac{\color{royalblue}{(x^2-4)'}(x)-(x^2-4)\color{royalblue}{(x)'}}{x^2}\]

zepdrix (zepdrix):

\[\Large =\frac{\color{royalblue}{(2x)}(x)-(x^2-4)\color{royalblue}{(1)}}{x^2}\]

zepdrix (zepdrix):

Mmm ok ok ok looks like you're on the right track. reduced? Hmm

OpenStudy (anonymous):

yay:) & yes because can't you simplify after you foil??

zepdrix (zepdrix):

No, after combining the terms in the numerator, you should probably leave it alone :) \[\Large \frac{x^2+4}{x^2}\] Hmm there is no foil :D See how the (x^2-4) is just being multiplied by 1? +_+

OpenStudy (anonymous):

Oh so you don't multiply the 2x and the x to get 2x^2?? :O

zepdrix (zepdrix):

You do ^^ sorry I skipped that step D:

OpenStudy (anonymous):

that's how I got (2x^2-x^2+4)/x^2 and then I assume you can reduce & get 2-1+4 = 5?? or am I wrong?

zepdrix (zepdrix):

You can combine the `x^2`s since they're like terms. \[\Large 2x^2-x^2 \quad=\quad x^2\]

zepdrix (zepdrix):

I'm not sure what you're doing with the 4 there though 0_o hmm

zepdrix (zepdrix):

Oh oh oh i see, you were trying to divide x^2 out?

zepdrix (zepdrix):

You forgot to divide x^2 from the last value, 4. It would leave you with 4/x^2, not just 4.

zepdrix (zepdrix):

That's why we want to leave it alone :O no division.

OpenStudy (anonymous):

yes but idk if that's right:o

OpenStudy (anonymous):

but wait doesn't the x^2 get cancelled out when it's reduced??

zepdrix (zepdrix):

\[\Large \frac{2x^2-x^2+4}{x^2} \quad\color{red}{\ne}\quad 2-1+4\] \[\Large \frac{2x^2-x^2+4}{x^2} \quad=\quad 2-1+\frac{4}{x^2}\]

OpenStudy (anonymous):

ohh :0 so it would be 1+4/x^2

OpenStudy (anonymous):

that's so true I forgot to divide that last part!

zepdrix (zepdrix):

Yes, you could write it like that if you want :) Or you can leave the x^2s on top which looks a little nicer. \(\Large \dfrac{x^2+4}{x^2}\qquad\)or\(\Large \qquad1+\dfrac{4}{x^2}\) Either way is fine ^^

OpenStudy (anonymous):

Got it :) thanks @zepdrix !!! :D

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

^.^